OpenStudy (anonymous):

Someone please help I have no idea how to approach this!!! The following polynomial represents a profit function for a certain production line, where x is a number of produced units, find a. the zeros and the multiplicity of each b. where the graph crosses or touches the x‐axis c. number of turning points d. explain the meaning of the above, if any f(x)=(x- 1/3)^2 〖(x-2)〗^3

OpenStudy (anonymous):

I have no idea how to this problem at all any and all help is appreciated and if you do help could you show your work so I can have it for notes for my test, thanks in advance.

OpenStudy (mathmale):

I'm going to assume that the polynomial in question is \[f(x)=(x-\frac{ 1 }{ 3 })^2\]

OpenStudy (anonymous):

okay, what would the second half represent/ and how do I proceed to a - d in regards to the polynomial?

OpenStudy (mathmale):

sorry for the interruption in our conversation. I had technical problems with this site at my end. I have no idea how "the first part" and 'the second part' are related; could you possibly share a screen shot of the original problem statement with me?

OpenStudy (mathmale):

Focusing only upon \[f(x)=(x-\frac{ 1 }{ 3 })^2,\]...and wanting to find the zero(s), I set this function = to zero:\[f(x)=(x-\frac{ 1 }{ 3 })^2=0, \] and then solve for x. How many roots would you suppose this equation has?

OpenStudy (anonymous):

2 2?

OpenStudy (mathmale):

\[f(x)=(x-\frac{ 1 }{ 3 })^2=0\] can be solved for x by taking the square root of both sides of this equation. Could you do that, please? type in your result.